OpenStudy (anonymous):

Is the inverse of y=1/x-5.... y^-1=1/x+5?

hartnn (hartnn):

no..

OpenStudy (anonymous):

Is it y^-1=x-5? @hartnn

hartnn (hartnn):

you need inverse function of y=1/(x-5) right ? try to isolate x from there

OpenStudy (anonymous):

Yeah if i isolate the x would it be y^-1= -1/x+5?

hartnn (hartnn):

oh, you mean \(\large \dfrac{1}{x}+5\) ? yes, thats the correct inverse function :)

OpenStudy (anonymous):

Thank you :)

hartnn (hartnn):

welcome ^_^